# Stochastic Path Planning in Correlated Obstacle Fields

### v1

@@ 1 Introduction
To tackle these challenges, we propose a two-stage policy learning strategy with the following key contributions: (i) Exploration guided by information gain: we augment the learning process with information bonus built upon mutual information, encouraging deeper exploration of uncertain and informative regions compared to traditional methods, with convergence property established. (ii) Flexible value function estimate approaches: in additional to classic Monte Carlo methods providing point estimates, our framework supports distribution RL to learn full cost distributions, providing stronger uncertainty quantification and more robust policies under high noise and risk. (iii) Offline-online integration: we develop a two-stage strategy enabling online adaptation, which outperforms standard rollout methods relying on static base policies. It consists of an offline phase that learns a high quality base policy via hybrid RL, and an online rollout phase with periodic base updates via a Bayesian mechanism that improves point estimates and distribution approximation. (iv) Belief updates with space reduction: blockage probabilities are refined using a Bayesian updating framework based on Gaussian Random Field model, with a search space reduction step guided by the posteriors to enhance computational efficiency.
@@ 4.1 Prior Information and Assumptions
where $\mu$ is the prior mean vector and $K$ is the covariance matrix. We set $\mu=0$ which centers the prior log-odds for each obstacle at zero, corresponding to a prior median (and mean, in the absence of uncertainty) of $\rho_{i}=0.5$ for all $i$. This setup indicates that each obstacle is equally likely to be true or false a priori. The matrix $K$ contains elements $\{K_{ij}=k(x_{i},x_{j})\}$ determined a kernel function $k$ capturing the inter-dependency between obstacles at two locations $x_{i}$ and $x_{j}$. We use a squared exponential kernel in spatial distance to model correlation such that nearby obstacles tend to have similar log-odds values, thereby capturing the intuition that spatially close obstacles tend to have similar blockage status:
@@
where $\sigma_{f}$ is and $l$ are kernel parameters controlling the strength and decay rate of spatial correlations.
Given a sequence of sensor readings and disambiguation outcomes for each obstacle, denoted as $\tilde{\rho_{i}}^{1},\ldots,\tilde{\rho_{i}}^{t}$, we aim to refine the probability estimates $P(Z_{i}=1|\tilde{\rho_{i}}^{1},...,\tilde{\rho_{i}}^{t}),i=1,2,...,n$. This framework allows us to update belief over obstacles that has not been directly sensed due to limited sensing range, by using the spatial correlation captured in the Gaussian prior.
@@ 4.4 Benefits of Posterior Updating based on the Correlation Structure
Theorem 4.1.
Let $\pi^{*}_{\mathrm{cor}}$ and $\pi^{*}_{\mathrm{ind}}$ denote the optimal traversal policies under posterior distributions with and without correlation information, respectively. Then, under the same environment,
$\mathbf{E}\left[\sum_{t=0}^{T}\mathcal{C}(S_{t},d^{\pi^{*}_{\mathrm{cor}}}(S_{t}))\right]\;\leq\;\mathbf{E}\left[\sum_{t=0}^{T}\mathcal{C}(S_{t},d^{\pi^{*}_{\mathrm{ind}}}(S_{t}))\right].$
@@
Let $\Phi_{\mathrm{cor}}$ denote the random variable representing posterior distributions over obstacle statuses when correlation information is incorporated (i.e., under the full Gaussian Random Field model). Similarly, let $\Phi_{\mathrm{ind}}$ denote the random variable when the posterior distribution is obtained under the assumption that obstacle statuses are independent.
Define a transformation $\Gamma$ that removes the correlation structure from any realization $\phi_{\mathrm{cor}}$ of $\Phi_{\mathrm{cor}}$, yielding a corresponding independent posterior:
$\Phi_{\mathrm{ind}}=\Gamma(\Phi_{\mathrm{cor}}).$
In the sense of Blackwell informativeness (Blackwell,, 1953; de Oliveira,, 2018), $\Phi_{\mathrm{ind}}$ is a garbling of $\Phi_{\mathrm{cor}}$, meaning that $\Phi_{\mathrm{cor}}$ is at least as informative as $\Phi_{\mathrm{ind}}$.
For any policy $\pi$ that maps belief states to decisions, let $\mathcal{C}(S_{t},d^{\pi}(S_{t}))$ denote the cost incurred at time $t$ under state $S_{t}$ and decision $d^{\pi}(S_{t})$. Then, by Blackwell’s Theorem, the optimal expected cost under the more informative signal (posterior with correlation) is no greater than that under the less informative signal:
$\min_{\pi\in\Pi_{\Phi_{\mathrm{cor}}}}\mathbf{E}\left[\sum_{t=0}^{T}\mathcal{C}\bigl(S_{t},d^{\pi}(S_{t})\bigr)\right]\;\leq\;\min_{\pi\in\Pi_{\Phi_{\mathrm{ind}}}}\mathbf{E}\left[\sum_{t=0}^{T}\mathcal{C}\bigl(S_{t},d^{\pi}(S_{t})\bigr)\right],$
where $\Pi_{\Phi_{\mathrm{cor}}}$ and $\Pi_{\Phi_{\mathrm{ind}}}$ denote the respective sets of feasible policies conditioned on their posterior models.
By definition of the optimal policies $\pi^{*}_{\mathrm{cor}}$ and $\pi^{*}_{\mathrm{ind}}$, the inequality becomes
$\mathbf{E}\left[\sum_{t=0}^{T}\mathcal{C}\bigl(S_{t},d^{\pi^{*}_{\mathrm{cor}}}(S_{t})\bigr)\right]\;\leq\;\mathbf{E}\left[\sum_{t=0}^{T}\mathcal{C}\bigl(S_{t},d^{\pi^{*}_{\mathrm{ind}}}(S_{t})\bigr)\right],$
which completes the proof. ∎
@@
This result implies that correlation-aware updating provides theoretical guarantees for improved performance. We also examine how additional information monotonically improves decision quality, further supporting the goal of uncertainty reduction in our posterior beliefs by sequentially gathering new observations through sensing or disambiguation.
@@
Let $X^{U}$ be the set of currently ambiguous obstacles and let $X^{U}_{1}\subseteq X^{U}_{2}\subseteq X^{U}$ be two observation sets, and $\Phi_{1}$, $\Phi_{2}$ denote the random variables representing the posterior distributions following observing those two sets, respectively. Let $\Pi_{1}$ and $\Pi_{2}$ be the sets of feasible policies with respect to their posterior models. Then the minimum expected total cost is monotonically decreasing in additional observations:
$\mathbf{E}\left[\mathcal{C}(X^{U}_{1})\right]\geq\mathbf{E}\left[\mathcal{C}(X^{U}_{2})\right],$
where $\mathbf{E}\left[\mathcal{C}(X)\right]:=\min_{\pi\in\Pi_{X}}\mathbf{E}\!\left[\sum_{t=0}^{T}\mathcal{C}\bigl(S_{t},d^{\pi}(S_{t})\bigr)\right].$ Equivalently, for the cost reduction function defined as $f(X):=\mathbf{E}\left[\mathcal{C}(\emptyset)\right]-\mathbf{E}\left[\mathcal{C}(X)\right]$, we have that $f$ is monotonically increasing: $f(X^{U}_{2})\geq f(X^{U}_{1})$.
@@
Let $\tilde{Y}_{1}$ and $\tilde{Y}_{2}$ denote the random log-odds vectors from sets $X^{U}_{1}$ and $X^{U}_{2}$ respectively, where $\tilde{Y}_{2}=\left(\tilde{Y}_{1},\tilde{Y}_{2\setminus 1}\right)$ and $\tilde{Y}_{2\setminus 1}$ represents the additional log-odds observations from $X^{U}_{2}\setminus X^{U}_{1}$.
For any realization of the observations:
$p(Y|\tilde{Y}_{1}=\tilde{\mathbf{y}}_{1})$ $=\int p(Y|\tilde{Y}_{1}=\tilde{\mathbf{y}}_{1},\tilde{Y}_{2\setminus 1}=\tilde{\mathbf{y}})\cdot p(\tilde{Y}_{2\setminus 1}=\tilde{\mathbf{y}}|\tilde{Y}_{1}=\tilde{\mathbf{y}}_{1})\,d\tilde{\mathbf{y}}$
$=\int p(Y|\tilde{Y}_{2}=(\tilde{\mathbf{y}}_{1},\tilde{\mathbf{y}}))\cdot p(\tilde{Y}_{2\setminus 1}=\tilde{\mathbf{y}}|\tilde{Y}_{1}=\tilde{\mathbf{y}}_{1})\,d\tilde{\mathbf{y}}$
This shows that for any realization, the posterior $\Phi_{1}$ is obtained by averaging the posterior $\Phi_{2}$ over the additional observations $\tilde{Y}_{2\setminus 1}$. Hence $\Phi_{1}$ is a garbling of $\Phi_{2}$.
By Blackwell’s Theorem (as in Theorem 4.1), the more informative signal $\Phi_{2}$ yields no greater minimum expected cost than $\Phi_{1}$, establishing the inequality $f(X^{U}_{2})\geq f(X^{U}_{1})$. ∎

### v2

@@ 1 Introduction
To tackle these challenges, we develop a statistical inference and computation pipeline that combines Bayesian inference with reinforcement learning (RL) and scalable search: (i) We derive GRF-based belief updates that refine posterior blockage probabilities from noisy, local signals and show how the resulting posteriors support principled search-space reduction (pruning) for planning. (ii) We propose a two-stage strategy: an offline optimistic policy-iteration (OPI) stage with an information-gain bonus to direct exploration toward informative regions, followed by an online rollout policy with periodic Bayesian updates to adapt as data accrue. (iii) Beyond Monte Carlo point estimates, we incorporate distributional RL to learn full cost distributions, improving uncertainty quantification and robustness under high noise/risk. (iv) We establish guarantees that correlation-aware beliefs weakly dominate their independence-coarsened counterparts in expected total cost, information gain is submodular under the linear–Gaussian sensing model, justifying greedy sensing policies, and the OPI/rollout scheme with posterior sampling converges under standard discounted or stochastic-shortest-path (SSP) conditions. Extensive simulations across sensing ranges, noise levels, and correlation regimes demonstrate consistent improvements over strong baselines.
@@ 3.1 Formulating SCOS as a Sequential Decision Problem
Based on the problem definition and sequential decision formulation, we work on a finite, undirected graph with bounded, nonnegative edge costs, and allow sensing, disambiguation at additional bounded costs. The goal state $g$ is absorbing since the traversal stops and no further cost incurs once $g$ is reached. Assuming there is at least one feasible path, possibly unnecessarily long, from $s$ to $g$, there exists a proper policy (i.e., one that reaches g almost surely from every state after finitely many steps). Consequently, our objective is undiscounted with a finite horizon. Throughout, we restrict the analysis to proper policies.
@@ 4.1 Prior Information and Assumptions
where $\mu$ is the prior mean vector and $K$ is the covariance matrix. We set $\mu=0$ which centers the prior log-odds for each obstacle at zero, corresponding to a prior median (and mean, in the absence of uncertainty) of $\rho_{i}=0.5$ for all $i$. This setup indicates that each obstacle is equally likely to be true or false a priori. The matrix $K$ contains elements $\{K_{ij}=k(x_{i},x_{j})\}$ determined by a kernel function $k$ capturing the inter-dependency between obstacles at two locations $x_{i}$ and $x_{j}$. We use a squared exponential kernel in spatial distance to model correlation such that nearby obstacles tend to have similar log-odds values, thereby capturing the intuition that spatially close obstacles tend to have similar blockage status:
@@
where $\sigma_{f}$ and $l$ are kernel parameters controlling the strength and decay rate of spatial correlations.
Given a sequence of sensor readings and disambiguation outcomes for each obstacle, denoted as $\tilde{\rho_{i}}^{1},\ldots,\tilde{\rho_{i}}^{t}$, we aim to refine the probability estimates $P(Z_{i}=1|\tilde{\rho_{i}}^{1},...,\tilde{\rho_{i}}^{t}),i=1,2,...,n$. This framework allows us to update belief over obstacles that have not been directly sensed due to limited sensing range, by using the spatial correlation captured in the Gaussian prior.
@@ 4.4 Benefits of Posterior Updating based on the Correlation Structure
4.4.1 Result on Correlation-Aware Dominance
We consider either a finite-horizon problem with $t\in\{0,\dots,T\},\;T<\infty$ on an underlying partially observed environment or an SSP (stochastic shortest path) setting with absorbing goal $g$. Single-stage costs are bounded: $\sup_{s,d}\mathcal{C}(s,d)<\infty$. At time $t$, the state is $S_{t}=\{V_{t},B_{t}\}$ with location $V_{t}$ and belief $B_{t}$. Let $b^{\mathrm{cor}}_{t}\in\mathcal{B}_{\mathrm{cor}}$ denote the correlation-aware belief (full joint posterior under the GRF model) and $U(b^{\mathrm{cor}}_{t},d_{t},Y_{t+1})$ the Bayesian update after action $d_{t}$ and observation $Y_{t+1}$. Policies are sequences $(\pi_{t})_{t}$ with measurable selectors $\pi_{t}:\,(V_{t},b)\mapsto d_{t}\in\mathcal{D}(S_{t})$ adapted to the canonical filtration.
We now provide the standing assumptions more explicitly:
  1. (A1)
    (Finite horizon and bounded costs) $T<\infty$ and $\sup_{s,d}\mathcal{C}(s,d)<\infty$.
  2. (A2)
    (Correlation-aware beliefs) $b^{\mathrm{cor}}_{t+1}=U(b^{\mathrm{cor}}_{t},d_{t},Y_{t+1})$ for all $t$.
  3. (A3)
    (Per-stage belief coarsening) There exists a measurable “coarsening” map $\Gamma:\mathcal{B}_{\mathrm{cor}}\to\mathcal{B}_{\mathrm{ind}}$ such that, pathwise,
    $b^{\mathrm{ind}}_{t}=\Gamma\!\big(b^{\mathrm{cor}}_{t}\big),\qquad b^{\mathrm{ind}}_{t+1}=\Gamma\!\Big(U\big(b^{\mathrm{cor}}_{t},d_{t},Y_{t+1}\big)\Big).$
    (For example, $\Gamma$ may replace the joint posterior by the product of its marginals.)
Theorem 4.1 (Correlation-Aware Dominance).
Under (A1)(A3), let
$J^{\star}_{\mathrm{cor}}:=\inf_{\pi}\,\mathbb{E}\!\Big[\textstyle\sum_{t}\mathcal{C}\big(S_{t},\pi_{t}(V_{t},b^{\mathrm{cor}}_{t})\big)\Big],\qquad J^{\star}_{\mathrm{ind}}:=\inf_{\tilde{\pi}}\,\mathbb{E}\!\Big[\textstyle\sum_{t}\mathcal{C}\big(S_{t},\tilde{\pi}_{t}(V_{t},b^{\mathrm{ind}}_{t})\big)\Big],$
be the optimal expected total costs when decisions are based on correlation-aware vs. coarsened (independent) beliefs, respectively. Then $J^{\star}_{\mathrm{cor}}\leq J^{\star}_{\mathrm{ind}}$.
@@ 4.4.1 Result on Correlation-Aware Dominance
Fix any admissible $\tilde{\pi}=(\tilde{\pi}_{t})$ on $(V_{t},b^{\mathrm{ind}}_{t})$ and define the replicated policy on correlated beliefs by composition,
$\pi_{t}(V_{t},b)\ :=\ \tilde{\pi}_{t}\!\big(V_{t},\Gamma(b)\big).$
By the stagewise coarsening property (i.e. Assumption (A3)), for every $t$ we have $\pi_{t}(V_{t},b^{\mathrm{cor}}_{t})=\tilde{\pi}_{t}(V_{t},b^{\mathrm{ind}}_{t})$, so the two policies induce identical action sequences and thus the same state/cost trajectories on every sample path. Therefore
$\mathbb{E}\!\Big[\textstyle\sum_{t}\mathcal{C}\big(S_{t},\pi_{t}(V_{t},b^{\mathrm{cor}}_{t})\big)\Big]=\mathbb{E}\!\Big[\textstyle\sum_{t}\mathcal{C}\big(S_{t},\tilde{\pi}_{t}(V_{t},b^{\mathrm{ind}}_{t})\big)\Big].$
Taking $\inf_{\tilde{\pi}}$ on the right yields $J^{\star}_{\mathrm{cor}}\leq J^{\star}_{\mathrm{ind}}$. That is, taking the infimum over $\tilde{\pi}$ on the right shows that the correlation-aware decision maker can mimic any independent-belief policy by first coarsening and then acting; therefore $J^{\star}_{\mathrm{cor}}\leq J^{\star}_{\mathrm{ind}}$. ∎
Equivalently, the theorem implies an inequality (or dominance) in expectations as well. Because the finite-horizon problem with finite state and action sets admits optimal policies, let
$\pi^{\star}_{\mathrm{cor}}\in\arg\min_{\pi}\,\mathbb{E}\!\Big[\textstyle\sum_{t=0}^{T}\mathcal{C}\big(S_{t},\pi_{t}(V_{t},b^{\mathrm{cor}}_{t})\big)\Big],\quad\pi^{\star}_{\mathrm{ind}}\in\arg\min_{\tilde{\pi}}\,\mathbb{E}\!\Big[\textstyle\sum_{t=0}^{T}\mathcal{C}\big(S_{t},\tilde{\pi}_{t}(V_{t},b^{\mathrm{ind}}_{t})\big)\Big].$
Then
$\mathbb{E}\!\Big[\textstyle\sum_{t=0}^{T}\mathcal{C}\big(S_{t},\pi^{\star}_{\mathrm{cor},t}(V_{t},b^{\mathrm{cor}}_{t})\big)\Big]\ \leq\ \mathbb{E}\!\Big[\textstyle\sum_{t=0}^{T}\mathcal{C}\big(S_{t},\pi^{\star}_{\mathrm{ind},t}(V_{t},b^{\mathrm{ind}}_{t})\big)\Big].$
Remark 4.2.
(i) On Blackwell Dominance. Suppose that for each stage $t$ there exists a Markov kernel $K_{t}$ such that, conditional on the latent state and the past history, the independent signal $Z^{\mathrm{ind}}_{t}$ is obtained by garbling the correlated signal $Z^{\mathrm{cor}}_{t}$:
$\mathcal{L}\!\big(Z^{\mathrm{ind}}_{t}\mid\text{state, history}\big)\;=\;\int K_{t}(\cdot\mid z)\,\mathcal{L}\!\big(Z^{\mathrm{cor}}_{t}\in dz\mid\text{state, history}\big),$
with $K_{t}$ independent of the latent state given $Z^{\mathrm{cor}}_{t}$. Then the correlated signal is stagewise more informative in Blackwell sense, and a standard dynamic extension implies the same value inequality as Theorem 4.1. In practice, verifying (A3) via a concrete belief coarsening map $\Gamma$ (e.g., product-of-marginals) is sufficient for Theorem 4.1 and avoids constructing signal-level garblings; note that such a $\Gamma$ need not correspond to a Blackwell garbling of raw signals.
(ii) Bellman Perspective. Let $V^{\mathrm{cor}}_{t}:\mathcal{B}_{\mathrm{cor}}\to\mathbb{R}$ and $V^{\mathrm{ind}}_{t}:\mathcal{B}_{\mathrm{ind}}\to\mathbb{R}$ denote the optimal cost-to-go at time $t$ in the correlation-aware and independent-belief problems, respectively. Define $\widetilde{V}_{t}:\mathcal{B}_{\mathrm{cor}}\to\mathbb{R}$ by $\widetilde{V}_{t}(b):=V^{\mathrm{ind}}_{t}\big(\Gamma(b)\big)$. Under (A3), a backward-induction argument shows
$V^{\mathrm{cor}}_{t}(b)\ \leq\ \widetilde{V}_{t}(b)\ =\ V^{\mathrm{ind}}_{t}\!\big(\Gamma(b)\big)\qquad\text{for all }b\in\mathcal{B}_{\mathrm{cor}},\ t=T,\dots,0,$
i.e., the dynamic programming operator is monotone under information coarsening after composing with $\Gamma$. Evaluating at $t=0$ yields Theorem 4.1.
Sketch of Proof. At $t=T$, the inequality is trivial. Assume it holds at $t+1$ and compare the Bellman minima for $V^{\mathrm{cor}}_{t}$ and $\widetilde{V}_{t}$; use that $b^{\mathrm{ind}}_{t+1}=\Gamma\!\big(U(b^{\mathrm{cor}}_{t},d,Y_{t+1})\big)$ (Assumption (A3)) to align the conditional expectations, and minimize over the same feasible action set $\mathcal{D}(S_{t})$.
@@
This result implies that correlation-aware updating provides theoretical guarantees for improved performance. We also examine how additional information monotonically improves decision quality through the following result, further supporting the goal of uncertainty reduction in our posterior beliefs by sequentially gathering new observations through sensing and disambiguation.
4.4.2 Result on Monotonicity with Additional Observations
Let $X^{U}$ denote the index set of ambiguous obstacles that could be observed (e.g., via sensing). That is, the set $X^{U}=\{x_{1},\ldots,x_{m}\}$ indexes all ambiguous obstacles whose statuses are uncertain and, in principle, observable (e.g., via sensing) at time 0. For any $X\subseteq X^{U}$, let $\widetilde{Y}_{X}:=\{\widetilde{Y}_{x}:\ x\in X\}$ be the the corresponding pre-decision observation vector collected from $X$ at time 0. Write $\mathcal{F}(X):=\sigma(\widetilde{Y}_{X})$ for the $\sigma$-field generated by those observations (together with the common prior). Admissible policies for $X$ are those that are measurable w.r.t. the canonical filtration generated by $\mathcal{F}(X)$ and the state/action history. Define the optimal expected total cost
$\mathbf{E}[\mathcal{C}(X)]\;:=\;\inf_{\pi\in\Pi_{X}}\;\mathbf{E}\!\Big[\textstyle\sum_{t=0}^{T}\mathcal{C}\big(S_{t},\ d_{t}\big)\Big],\quad\text{where }d_{t}=\pi_{t}(\text{history up to }t;\ \mathcal{F}(X)).$
Thus $X^{U}_{1}\subseteq X^{U}_{2}\subseteq X^{U}$ represents increasing observational richness:
$\mathcal{F}(X^{U}_{1})\ \subseteq\ \mathcal{F}(X^{U}_{2}),$
so a policy admissible under $\mathcal{F}(X^{U}_{1})$ is also admissible under $\mathcal{F}(X^{U}_{2})$. Intuitively, the decision maker with access to $X^{U}_{2}$ can condition on all information available under $X^{U}_{1}$ plus the additional signals for $X^{U}_{2}\setminus X^{U}_{1}$ (and can always ignore extra coordinates if desired).
@@ 4.4.2 Result on Monotonicity with Additional Observations
Let $X^{U}_{1}\subseteq X^{U}_{2}\subseteq X^{U}$. Then the optimal expected cost is monotone in the observation set:
$\mathbf{E}\big[\mathcal{C}(X^{U}_{2})\big]\ \leq\ \mathbf{E}\big[\mathcal{C}(X^{U}_{1})\big].$
Equivalently, for the cost-reduction function $f(X):=\mathbf{E}[\mathcal{C}(\emptyset)]-\mathbf{E}[\mathcal{C}(X)]$, we have $f(X^{U}_{2})\geq f(X^{U}_{1})$ (monotone nondecreasing).
@@
Since $X^{U}_{1}\subseteq X^{U}_{2}$, we have $\mathcal{F}(X^{U}_{1})\subseteq\mathcal{F}(X^{U}_{2})$. By the definition of admissible policies, any $\mathcal{F}(X^{U}_{1})$–adapted policy is also $\mathcal{F}(X^{U}_{2})$–adapted; hence $\Pi_{X^{U}_{1}}\subseteq\Pi_{X^{U}_{2}}$. Therefore
$\mathbf{E}[\mathcal{C}(X^{U}_{2})]=\inf_{\pi\in\Pi_{X^{U}_{2}}}\mathbf{E}\!\Big[\textstyle\sum_{t=0}^{T}\mathcal{C}(S_{t},d_{t})\Big]\ \leq\ \inf_{\pi\in\Pi_{X^{U}_{1}}}\mathbf{E}\!\Big[\textstyle\sum_{t=0}^{T}\mathcal{C}(S_{t},d_{t})\Big]=\mathbf{E}[\mathcal{C}(X^{U}_{1})],$
as claimed. ∎
Alternative Constructive Proof:.
Because $X^{U}_{1}\subseteq X^{U}_{2}$, we have $\mathcal{F}(X^{U}_{1})\subseteq\mathcal{F}(X^{U}_{2})$ and there is a measurable projection $\mathrm{proj}_{1}:\Omega\to$ realizations of $\mathcal{F}(X^{U}_{1})$ that discards the extra observations in $X^{U}_{2}\setminus X^{U}_{1}$. Take any admissible policy $\pi^{(1)}\in\Pi_{X^{U}_{1}}$.
Define the restriction map $r:\widetilde{Y}_{X^{U}_{2}}\to\widetilde{Y}_{X^{U}_{1}}$ that drops coordinates in $X^{U}_{2}\setminus X^{U}_{1}$. Given $\pi^{(1)}\in\Pi_{X^{U}_{1}}$, define $\pi^{(2)}\in\Pi_{X^{U}_{2}}$ by $\pi^{(2)}_{t}(\text{history},\widetilde{Y}_{X^{U}_{2}}):=\pi^{(1)}_{t}(\text{history},r(\widetilde{Y}_{X^{U}_{2}}))$. Then the two policies take identical actions pathwise, yielding the same cost; taking infima gives the result.
Remark 4.4 (When a Blackwell view applies).
Let the raw signals be $Z_{2}=(\widetilde{Y}_{X^{U}_{1}},\,\widetilde{Y}_{X^{U}_{2}\setminus X^{U}_{1}})$ and $Z_{1}=\widetilde{Y}_{X^{U}_{1}}$. Define the (deterministic) Markov kernel $K$ by $K(A\mid z_{2}):=\mathbf{1}\{\mathrm{proj}(z_{2})\in A\}$, where $\mathrm{proj}$ drops the coordinates in $X^{U}_{2}\setminus X^{U}_{1}$. Then, conditional on the latent state and past history,
$\mathcal{L}\!\big(Z_{1}\mid\text{state, history}\big)=\int K(\cdot\mid z_{2})\,\mathcal{L}\!\big(Z_{2}\in dz_{2}\mid\text{state, history}\big),$
so $Z_{2}$ Blackwell–dominates $Z_{1}$ in the one-shot sense. Consequently, any Bayes decision problem based on these signals has value (risk) under $Z_{2}$ no worse than under $Z_{1}$, and the monotonicity conclusion follows. Our proof in the main text circumvents signal-level conditions by working directly with $\sigma$-fields and sequential policy classes.
Remark 4.5 (Sequential variants).
If observations are acquired across time and the richer filtration $\{\mathcal{F}_{t}(X^{U}_{2})\}_{t=0}^{T}$ satisfies $\mathcal{F}_{t}(X^{U}_{1})\subseteq\mathcal{F}_{t}(X^{U}_{2})$ for every $t$, then the admissible policy classes obey $\Pi_{X^{U}_{1}}\subseteq\Pi_{X^{U}_{2}}$. Hence the same replication / policy-class inclusion argument yields $\mathbf{E}[\mathcal{C}(X^{U}_{2})]\leq\mathbf{E}[\mathcal{C}(X^{U}_{1})]$ for the finite-horizon problem (and analogously for SSP with the sum taken to the hitting time).